# TIGERScore: Towards Building Explainable Metric for All Text Generation Tasks

### v3

@@ Abstract
We present TIGERScore, a Trained metric that follows Instruction Guidance to perform Explainable, and Reference-free evaluation over a wide spectrum of text generation tasks. Different from other automatic evaluation methods that only provide arcane scores, TIGERScore is guided by natural language instruction to provide error analysis to pinpoint the mistakes in the generated text. Our metric is based on LLaMA-2, trained on our meticulously curated instruction-tuning dataset MetricInstruct which covers 6 text generation tasks and 23 text generation datasets. The dataset consists of 42K quadruple in the form of (instruction, input, system output $\rightarrow$ error analysis). We collected the ‘system outputs’ through from a large variety of models to cover different types of errors. To quantitatively assess our metric, we evaluate its correlation with human ratings on 5 held-in datasets, 2 held-out datasets and show that TIGERScore can achieve the open-source SoTA correlation with human ratings across these datasets and almost approaches GPT-4 evaluator. As a reference-free metric, its correlation can even surpass the best existing reference-based metrics. To further qualitatively assess the rationale generated by our metric, we conduct human evaluation on the generated explanations and found that the explanations are 70.8% accurate. Through these experimental results, we believe TIGERScore demonstrates the possibility of building universal explainable metrics to evaluate any text generation task.
@@ 1 Introduction
Evaluation of natural language generation tasks is a long-standing challenging problem. With the recent advancement of large pre-trained language models (Brown et al., 2020; OpenAI, 2023; Touvron et al., 2023), text generation models have become more widely adopted in real-world applications than ever. Newly developed text-generative models are being deployed across a wide range of downstream applications. As more and more people use these generative models, there is an increasing need to develop trustworthy evaluation metrics. Though GPT-4 has shown to achieve strong evaluation correlation with humans (Zheng et al., 2023), the open-sourced metrics are lagging significantly behind and mainly suffer from specific issues:
    Dependency on references: Some evaluation metrics like ROUGE (Lin, 2004), BLEU (Papineni et al., 2002b), COMET (Rei et al., 2020), InstructScore (Xu et al., 2023c) would require gold references to measure the quality. These metrics compare the generated output against one or more reference texts to assign the evaluation score. However, such assumption can be highly unrealistic in many downstream applications where the golden reference is hard to collect.
@@
    Lack of attribution:: Some evaluation metrics tend to directly output a score without any attributions, e.g., where the errors occur and why. For instance, BARTScore (Yuan et al., 2021), BERTScore (Zhang et al., 2019) and GPTScore (Fu et al., 2023) adopt the pre-trained language models’ log likelihood as the evaluation metric. Such metrics do not provide the location and reason for the assigned score, which limits their trustworthiness and reliability.
Figure 1: The upper part shows the input and output format of TIGERScore. The lower part shows the Kendall’s correlation of different metrics w.r.t human ratings, from which we can see TIGERScore achieves the SoTA correlation on most generation tasks.
To address these issues, we propose a novel metric, TIGERScore, a Trained metric that follows Instruction Guidance to perform Explainable and Reference-free evaluation. As shown in Figure 1, the input to TIGERScore consists of an instruction describing the task definition, the task input, and the system output. TIGERScore is capable of generating breakdown error analysis that can (1) locate each mistake, (2) explain the mistake and suggest revisions, and (3) provide a penalty score (between [-5, -0.5]) for each mistake. The final score can be calculated by summing all the penalty scores.
TIGERScore is built by fine-tuning LLMs on our curated MetricInstruct dataset, which contains a total of 42K examples of (instruction, input, system output, error analysis), obtained from 23 diverse text generation datasets. The dataset includes system outputs from more than 50 systems, covering a wide variety of errors. The error analysis is curated by prompting GPT-4 (OpenAI, 2023) and filtered through various strategies. The tuned model TIGERScore has shown the highest overall Kendall’s correlation with human ratings on seven major text generation tasks as depicted in Figure 1. TIGERScore is highly convenient to use because it does not require any additional reference. We evaluate TIGERScore on 5 held-in datasets and 2 held-out datasets. As a reference-free metric, TIGERScore surpasses the best reference-free metrics (GPTScore (Fu et al., 2023)) by 15% and surpasses the best reference-based metrics (COMET-22 (Rei et al., 2022a)) by a margin of 8% in terms of Kendall’s score. On the two held-out datasets, TIGERScore demonstrates unprecedented generalization capabilities. We further employ humans to evaluate the explanations generated by TIGERScore and found that over 70% of the generated explanations are highly accurate and trustworthy. Our analysis shows that the success of TIGERScore is attributed to three key aspects in MetricInstruct: (1) dataset diversity, (2) error coverage, and (3) high quality, which enable TIGERScore to generalize better than any other metric.
@@ 2.1 Background
The pursuit of improved metrics for text evaluation has been a significant focus since the inception of language models. Automatic n-gram-based metrics (Elliott & Keller, 2014; Callison-Burch et al., 2006; Isozaki et al., 2010) have always served as the default metric, computing the n-gram match F-1 score with the reference text until research highlighted their significant weaknesses in aligning with human preferences. Later, neural metrics were proposed to capture semantic similarity by either computing based on neural representation (Zhang et al., 2019; Yuan et al., 2021) or directly fine-tuning with human preferences (Rei et al., 2020). It has also been demonstrated that multi-aspect scores, using the logits of large language models with well-designed instructions as prompts (Fu et al., 2023), could achieve an even higher correlation.
@@ 2.2 Problem Formulation
$\{...,E_{i},...\}=\{...,(l_{i},a_{i},e_{i},s_{i}),...\}=f(I,x,y^{\prime})$ (1)
where $(l_{i},a_{i},e_{i},s_{i})$ denotes specific information of the error $E_{i}$. Specifically, $l_{i}$ points to the location of the error, and $a_{i}\in A_{T}$ is a pre-defined aspect to which this error belongs. $e_{i}$ comprises both the explanation of this error and its revision suggestion. $s_{i}$ is the penalty score reduced for this error, which lies in $[-5,-0.5]$. The final score of $y^{\prime}$ is computed as the sum of all the penalty scores: $s_{y^{\prime}}=\sum_{i}{s_{i}}$. The range of the final score lies within $(-\infty,0]$, where 0 means perfect generation and a lower score indicates worse quality, more and severer errors.
@@ 2.3 Multi-Aspect Evaluation
Task Aspect
Definition
Instruct Comprehension
Evaluates how well the output understands the given instruction.
Accuracy
Measures the correctness of the output in relation to the instruction and the paired input context.
Informativeness
Assesses the relevancy and usefulness of the information provided by the output.
Coherence
Evaluates how logically the output flows and connects.
@@ 2.4 Training Setup
TIGERScore is finetuned on Llama-2-7B and Llama-2-13B (Touvron et al., 2023) respectively with both’s batch size being 128. The prompt template used for both the fine-tuning and inference are given in  Table 9. The model’s maximum context length is set to 1024. We use a cosine learning scheduler with the warmpup ratio being 0.1 We finetune the 7B version on 4 A100 GPUs for 3 epochs with a learning rate of 2e-5. And 13B version is run on 8 A100 GPUs for 2 epochs with a learning rate of 1e-5. Inference of TIGERScore is conducted on a single A100 GPU with the assistance of the VLLM toolkit to increase the speed. (Kwon et al., 2023)
@@ 3.1 Diverse Dataset Source
In addition, we meticulously selected datasets for each task to ensure diverse coverage across the knowledge domain, as shown in  Table 2. For instance, in the case of LF-QA, we utilize ASQA (Stelmakh et al., 2022b) to address ambiguous factoid questions, while FeTaQA (Nan et al., 2022b) is employed to handle challenges related to tabular source question answering. We then selected samples from each dataset’s training set, following predefined constraints like maximum input and output lengths, and topics balancing of instances and so on.
@@ 3.2 Broad Coverage of Errors
As shown in  Figure 2, our “system outputs” come from two channels, namely real-world system outputs and synthetic incorrect outputs, representing 2 main components of MetricInstruct
As illustrated in  Figure 2, for those self-generated outputs, we sample 5 different outputs using top-p sampling for each instance. After that, we use BARTScore to sorted the 5 outputs. Outputs with lower BARTScore is heuristically considered to contain more errors, and will be preferred to be chosen as the final hypothesis output $y^{\prime}$ to be evaluated.
Subsequently, we utilized meticulously designed prompting templates (see in  A.4) to elicit standardized error analysis from GPT-4. The main idea is to provide them with the instruction $I$, input $x$, the reference output $y$, and system output $y^{\prime}$ along with the definitions of pre-defined aspects $A_{T}$, to query the OpenAI models to generate the list of errors, as described in subsection 2.2. We also report their correlation performance in Table 4, showcasing the reference-based ChatGPT results.
@@
While real-world system outputs ensure the error distribution is aligned with practical scenarios, they might be overly tailored to bias of these limited systems and omit certain error cases their outputs did not cover. Therefore, we also synthesize incorrect outputs that can balance under-represented errors by prompting GPT-4 with templates.
@@
Secondly, to improve TIGERScore’s generalization ability to the unseen tasks, we also prompt GPT-4 to generate data with customized error aspects tailored to individual instructions, using the template in  Table 18. We hypothesize that the manually designed aspects, while logical, may not be specific enough for tasks like instruction-following where instructions vary greatly. 10k data are sampled from Alpaca-52k (Taori et al., 2023) for this generation part.
@@ 4 Experiments
Summarization SummEval 16 Systems 100 1600
Translation WMT-22 (zh-en) 18 Systems 1875 33750
Data to Text WebNLG-2020 18 Systems 179 2848
Long-form QA ASQA, FeTaQA, CosmosQA, ELI5 4 Systems 400 1600
@@
Instruction-Following LIMA,AlpacaEval 9 Systems 500 4500
Story Generation OpenMEVA (ROC) 5 Systems 200 1000
Table 3: Overview of our main evaluation datasets. “# Samples” is computed by multipling “# Inputs” with the number of outputs systems, representing the total instances that a metric need to compute. Performance of baselines and TIGERScore on these datasets are reported in  Table 4.
4.1 Evaluation Datasets
Human preference scores are necessary to conduct the correlation analysis. Those datasets with official system outputs released are usually accompanied by systematic human preference scores, like the WMT-MQM score for the translation task. However, these scores are not available for tasks like long-form QA, MathQA, and instruction-following.
4.2 Baselines
Reference-based: We choose popular metrics, including BLEU (Papineni et al., 2002a), ROUGE (Lin, 2004), BERTScore (Zhang et al., 2019), BLEURT (Sellam et al., 2020) and BARTScore (Yuan et al., 2021). Recent emerging metrics are also included, like COMET-22 (Rei et al., 2022a),UniEval (Zhong et al., 2022b), GPTScore (Fu et al., 2023), and InstructScore (Xu et al., 2023c). Specifically, we use BARTScore-ref to denote that we adopt the ref-hypo scoring type. For GPTScore, we use FLAN-T5-base (Chung et al., 2022) as base models and use GPTScore-ref to denote that we adopt the f-1 average of the ref-hypo and hypo-ref scores. We also report the few-shot results of GPT-3.5-turbo with the same prompting templates as those used for generating real-world-channel data. Reporting this baseline will help us understand whether TIGERScore has surpassed its easy substitute using the cheap OpenAI model.
Reference-free: We choose BARTScore, GPTScore and COMETKiwi (Rei et al., 2022b) as reference-free baselines to compare. Specifically, we use BARTScore-src to denote the src-hypo scoring type, thus making it a reference-free metric. For GPTScore, we still use the FLAN-T5-base model and use GPTScore-src to denote that we use the src-hypo scoring type. We also include the frequently used 0-shot results of Llama-2-13b (Touvron et al., 2023) and GPT-4 (OpenAI, 2023)for better comparison and to know more about the performance gaps with GPT-4.
@@ 4.3 Main Results
We present a comprehensive analysis of TIGERScore across all 6 held-in tasks and 1 held-out task in  Table 4 reporting Kendall correlation results. Additionally, we provide supplementary results on Pearson and Spearman correlations in the Appendix (see Table 25 and Table 26). We average the performance across these tasks to gauge the general ability of the model.
Our results highlight the significant advantages of TIGERScore over other reference-free metrics. Notably, TIGERScore has surpassed all other reference-free metrics in Kendall correlation. In Pearson correlation, TIGERScore is the highest for 6 out of 7 tasks. This underscores the robustness and consistency of TIGERScore in evaluating text generation tasks.
Compared with reference-based baselines, TIGERScore generally outperforms most reference-based metrics. However, it does score lower than some task-specific metrics like UniEval (summ) in summarization, COMET-22 in translation. We consider these discrepancies acceptable, because the compared metrics are reference-based and all specifically fine-tuned for a single task. Remarkably, in four other tasks, TIGERScore exceeds the correlation scores of all reference-based metrics, in both single-task and all-task average aspects, reaffirming its effectiveness as a universal metric.
Furthermore, TIGERScore achieves significantly higher overall correlation, compared with the cheap API substitute GPT-3.5-Turbo (few-shot) and the Llama-2-13b-chat (0-shot), proving its effectiveness. What’s exciting to note is that TIGERScore-13b can achieve comparable correlation performance with GPT-4 (0-shot), even higher on some tasks like summarization, translation, data2text, and story generation. The average scores across all 7 tasks are also approaching GPT-4.

### v4

@@ Abstract
We present TIGERScore [^1] Project website: https://tiger-ai-lab.github.io/TIGERScore/, a Trained metric that follows Instruction Guidance to perform Explainable, and Reference-free evaluation over a wide spectrum of text generation tasks. Different from other automatic evaluation methods that only provide arcane scores, TIGERScore is guided by natural language instruction to provide error analysis to pinpoint the mistakes in the generated text. Our metric is based on LLaMA-2, trained on our meticulously curated instruction-tuning dataset MetricInstruct which covers 6 text generation tasks and 23 text generation datasets. The dataset consists of 42K quadruple in the form of (instruction, input, system output $\rightarrow$ error analysis). We collected the ‘system outputs’ through from a large variety of models to cover different types of errors. To quantitatively assess our metric, we evaluate its correlation with human ratings on 5 held-in datasets, 2 held-out datasets and show that TIGERScore can achieve the open-source SoTA correlation with human ratings across these datasets and almost approaches GPT-4 evaluator. As a reference-free metric, its correlation can even surpass the best existing reference-based metrics. To further qualitatively assess the rationale generated by our metric, we conduct human evaluation on the generated explanations and found that the explanations are 70.8% accurate. Through these experimental results, we believe TIGERScore demonstrates the possibility of building universal explainable metrics to evaluate any text generation task.
@@ 1 Introduction
Figure 1: The upper part shows the input and output format of TIGERScore, which takes an instruction and an input context, along with the model-generated output that is to be evaluated as the inputs, and output detailed breakdown error analysis with explanations. It is finetuned on our curated dataset MetricInstruct. The lower part shows the Kendall’s correlation of different metrics w.r.t human ratings, from which we can see TIGERScore achieves the SoTA correlation on most generation tasks.
Evaluation for natural language generation tasks is a long-standing challenging problem. With the recent advancement of large pre-trained language models (Brown et al., 2020; OpenAI, 2023; Touvron et al., 2023), text generation models have become more widely adopted in real-world applications than ever. Newly developed text-generative models are being deployed across a wide range of downstream applications. As more and more people use these generative models, there is an increasing need to develop trustworthy evaluation metrics. Though GPT-4 has shown to achieve strong evaluation correlation with humans (Zheng et al., 2023), the open-sourced metrics are lagging significantly behind and mainly suffer from specific issues:
    Dependency on references: Some evaluation metrics like ROUGE (Lin, 2004), BLEU (Papineni et al., 2002b), COMET (Rei et al., 2020), InstructScore (Xu et al., 2023c) would require gold references to measure the quality. These metrics compare the generated output against one or more reference texts to assign the evaluation score. However, such an assumption can be highly unrealistic in many downstream applications where the gold reference is hard to collect.
@@
    Lack of attribution: Some evaluation metrics tend to directly output a score without any attributions, e.g., where the errors occur and why. For instance, BARTScore (Yuan et al., 2021), BERTScore (Zhang et al., 2019) and GPTScore (Fu et al., 2023) adopt the pre-trained language models’ log likelihood as the evaluation metric. Such metrics do not provide the location and reason for the assigned score, which limits their trustworthiness and reliability.
To address these issues, we propose a novel metric, TIGERScore, a Trained metric that follows Instruction Guidance to perform Explainable and Reference-free evaluation. As shown in Figure 1, the input to TIGERScore consists of an instruction describing the task definition, the task input, and the system output. TIGERScore is capable of generating breakdown error analysis that can (1) locate each mistake, (2) explain the mistake and suggest revisions, and (3) provide a penalty score (between [-5, -0.5]) for each mistake. The final score for the system output can be calculated by summing up all the penalty scores.
TIGERScore is built by fine-tuning LLMs on our curated MetricInstruct dataset, which contains a total of 42K examples of (instruction, input, system output, error analysis), obtained from 23 diverse text generation datasets. The dataset includes system outputs from more than 50 systems, covering a wide variety of errors. The error analysis is curated by prompting GPT-4 (OpenAI, 2023) and filtered through various strategies. The tuned model TIGERScore has shown the highest overall Kendall’s correlation with human ratings on seven major text generation tasks as depicted in Figure 1. TIGERScore is highly convenient to use because it does not require any additional reference. We evaluate TIGERScore on 5 held-in datasets and 2 held-out datasets. As a reference-free metric, TIGERScore surpasses the best reference-free metrics (GPTScore (Fu et al., 2023)) by 15% and surpasses the best reference-based metrics (COMET-22 (Rei et al., 2022a)) by a margin of 8% in terms of Kendall’s score. On the two held-out datasets, TIGERScore demonstrates unprecedented generalization capabilities. We further employ humans to evaluate the explanations generated by TIGERScore and found that over 70% of the generated explanations are highly accurate and trustworthy. Our analysis shows that the success of TIGERScore is attributed to three key aspects in our curated MetricInstruct: (1) dataset diversity, (2) error coverage, and (3) high quality, which enable TIGERScore to generalize better on unseen tasks than any other metric.
@@ 2.1 Background
The pursuit of improved metrics for text evaluation has been a significant focus since the inception of language models. Automatic n-gram-based metrics (Elliott & Keller, 2014; Callison-Burch et al., 2006; Isozaki et al., 2010) have always served as the default metric, computing the n-gram match F-1 score with the reference text until research highlighted their significant weaknesses in aligning with human preferences. Later, neural metrics were proposed to better capture the semantic similarity in addition to mere morphological similarity only by either computing based on neural representation (Zhang et al., 2019; Yuan et al., 2021) or directly fine-tuning with human preferences (Rei et al., 2020). It has also been demonstrated that multi-aspect scores, using the logits of large language models with well-designed instructions for various aspects as prompts (Fu et al., 2023), could achieve an even higher correlation.
@@ 2.2 Problem Formulation
$\small\{...,E_{i},...\}=\{...,(l_{i},a_{i},e_{i},s_{i}),...\}=f(I,x,y^{\prime})$ (1)
where $(l_{i},a_{i},e_{i},s_{i})$ denotes specific information of the error $E_{i}$. Specifically, $l_{i}$ points to the location of the error, and $a_{i}\in A_{T}$ is a pre-defined aspect to which this error belongs. $e_{i}$ comprises both the explanation of this error and its revision suggestion. $s_{i}$ is the penalty score reduced for this error, which lies in $[-5,-0.5]$. The final score of $y^{\prime}$ is computed as the sum of all the penalty scores: $s_{y^{\prime}}=\sum_{i}{s_{i}}$. The range of the final score lies within $(-\infty,0]$, where 0 means perfect generation and a lower score indicates worse quality and more severe errors. However, in practice, the number of detected errors is affected by both the training data and the generation length, thus leading to less than 10 errors most of the time during the evaluation.
Instruct
Aspect Definition
Comprehension Evaluates how well the output understands the given instruction.
Accuracy Measures the correctness of the output in relation to the instruction and the paired input context.
Informativeness Assesses the relevancy and usefulness of the information provided by the output.
Coherence Evaluates how logically the output flows and connects.
@@ 2.4 Training Setup
TIGERScore is finetuned on Llama-2-7B and Llama-2-13B (Touvron et al., 2023) respectively with both’s batch size being 128. The prompt templates used for both the fine-tuning and inference are given in  Table 9. The model’s maximum context length is set to 1024. We use a cosine learning scheduler with the warmpup ratio being 0.1 We finetune the 7B version on 4 A100 GPUs (80GB) for 3 epochs with a learning rate of 2e-5. And 13B version is run on 8 A100 GPUs (80GB) for 2 epochs with a learning rate of 1e-5. Inference of TIGERScore is conducted on a single A100 GPU with the assistance of the VLLM toolkit to increase the speed. (Kwon et al., 2023)
@@ 3.1 Diverse Dataset Source
In addition, we meticulously selected datasets for each task to ensure diverse coverage across the knowledge domain, as shown in  Table 2. For instance, in the case of the LF-QA task, we utilize ASQA (Stelmakh et al., 2022b) to include knowledge about ambiguous factoid questions, while FeTaQA (Nan et al., 2022b) is employed to handle challenges related to tabular source question answering. We then selected samples from each dataset’s training set, following the predefined constraints like maximum input and reference output lengths, topics balancing of instances, and so on.
@@ 3.2 Broad Coverage of Errors
As shown in  Figure 2, our “system outputs” come from two channels, namely real-world system outputs and synthetic incorrect outputs, representing 2 main components of MetricInstruct.
As illustrated in  Figure 2, for those self-generated outputs, we sample 5 different outputs using top-p sampling for each instance using various output systems. After that, we use BARTScore to sort the 5 outputs. Outputs with lower BARTScore are heuristically considered to contain more errors and will be preferred to be chosen as the final hypothesis output $y^{\prime}$ to be evaluated.
Subsequently, we utilized meticulously designed prompting templates (see in  A.7) to elicit standardized error analysis from GPT-4. The main idea is to provide them with the instruction $I$, input $x$, the reference output $y$, and system output $y^{\prime}$ along with the definitions of pre-defined aspects $A_{T}$, to query the OpenAI models to generate the list of errors, as described in subsection 2.2. We also report their correlation performance in Table 4, showcasing the reference-based ChatGPT results.
@@
While real-world system outputs ensure the error distribution is aligned with practical scenarios, they might be overly tailored to the bias of these limited systems and omit certain error cases their outputs fail to cover. Therefore, we decided to synthesize more incorrect outputs that can balance under-represented errors by prompting GPT-4 with our designed templates.
@@
Secondly, to improve TIGERScore’s generalization capability in evaluation for tasks or errors it has not previously encountered, we employ a strategy that involves prompting GPT-4 to generate data with customized error aspects tailored to individual instructions. Our rationale is that while the manually designed aspects are logical, they may not be specific enough for tasks that involve following instructions where various types of instructions will be included. Besides, this can also prevent the trained model from being overfitted to the pre-defined aspects in  subsection A.6. To do this, we supplement the first part of synthetic data by sampling 10k data points from the Alpaca-52k dataset and using the template in  Table 19 to prompt GPT-4 to get this second part of synthetic data.
@@ 4 Experiments
4.1 Evaluation Datasets
Human preference scores are necessary to conduct the correlation analysis. To get the gold preference scores we used to conduct the evaluation experiments, we collected either existing human ratings like WMT-22 MQM scores for translation, or GPT-4 rated scores for Long-form QA. Details of the curation are shown in  subsection A.5.
4.2 Baselines
Reference-based:
We choose popular metrics, including BLEU (Papineni et al., 2002a), ROUGE (Lin, 2004), BERTScore (Zhang et al., 2019), BLEURT (Sellam et al., 2020) and BARTScore (Yuan et al., 2021). Recent emerging metrics are also included, like COMET-22 (Rei et al., 2022a),UniEval (Zhong et al., 2022b), GPTScore (Fu et al., 2023), and InstructScore (Xu et al., 2023c). Specifically, we use BARTScore-ref to denote that we adopt the ref-hypo scoring type. For GPTScore, we use FLAN-T5-base (Chung et al., 2022) as base models and use GPTScore-ref to denote that we adopt the f-1 average of the ref-hypo and hypo-ref scores. We also report the zero-shot results of GPT-3.5-turbo with the same prompting templates as those used for generating real-world channel data. Reporting this baseline will help us understand whether TIGERScore has surpassed its easy substitute using the cheap OpenAI model, thus proving its effectiveness.
Reference-free:
We choose BARTScore, GPTScore and COMETKiwi (Rei et al., 2022b) as reference-free baselines to compare. Specifically, we use BARTScore-src to denote the src-hypo scoring type, thus making it a reference-free metric. For GPTScore, we still use the FLAN-T5-base model and use GPTScore-src to denote that we use the src-hypo scoring type. We also include the frequently used 0-shot results of Llama-2-13b (Touvron et al., 2023) and GPT-4 (OpenAI, 2023)for better comparison and to know more about the performance gaps with GPT-4.
@@ 4.2 Baselines
Summ SummEval 16 Systems 100 1600
Trans WMT-22 (zh-en) 18 Systems 1875 33750
D2T WebNLG-2020 18 Systems 179 2848
LF-QA ASQA, FeTaQA, 4 Systems 400 1600
CosmosQA, ELI5
@@
Instruct LIMA,AlpacaEval 9 Systems 500 4500
Story-Gen OpenMEVA (ROC) 5 Systems 200 1000
Table 3: Overview of our main evaluation datasets. “# Samples” is computed by multipling “# Inputs” with the number of outputs systems, representing the total instances that a metric need to compute.
@@ 4.3 Main Results
We present a comprehensive analysis of TIGERScore across all 5 held-in tasks and 2 held-out task in  Table 4 reporting Kendall correlation results. Additionally, we provide supplementary results on Pearson and Spearman correlations in the Appendix (see Table 25 and Table 26). We average the performance across these tasks to gauge the general ability of the model.
Our results highlight the significant advantages of TIGERScore over other reference-free metrics. Notably, it has surpassed all other reference-free metrics in Kendall correlation. In Pearson correlation, it is the highest for 6 out of 7 tasks. This underscores the robustness and consistency of TIGERScore in evaluating text generation tasks.
Compared with reference-based baselines, TIGERScore generally outperforms most reference-based metrics. However, it does score lower than some task-specific metrics like UniEval (summ) in summarization, COMET-22 in translation. We consider these discrepancies acceptable, because the compared metrics are reference-based and all specifically fine-tuned for a single task. Furthermore, TIGERScore achieves significantly higher overall correlation, compared with the cheap API substitute GPT-3.5-Turbo (zero-shot) and the Llama-2-13b-chat (zero-shot), proving its effectiveness. What’s exciting to note is that TIGERScore-13b can achieve comparable correlation performance with GPT-4 (zero-shot), even higher on summarization, translation, data2text, and story generation.